OpenStudy (anonymous):

At the Boston Children's Museum, it was estimated 38% of all visitors are from in state. On Saturday, 8,000 people attended the museum. What is the number of out of state people in attendance? (

hero (hero):

is/of = %/100 x/8000 = 38/100 100x = 8000*38 x = (8000*38)/100 x = 80*38 x = 3040